# Training Optimal Large Diffusion Language Models

### v1

@@ Abstract
We introduce Quokka, the first systematic scaling law for diffusion language models (DLMs), encompassing both compute-constrained and data-constrained regimes, and studying the key modeling and optimization designs. Quokka is a good friend of Chinchilla and provides wider scopes. We hope the results would bring short-term practical guidance in DLMs training and long-term inspirations for the whole AI community. We summarize some takeaways below:
@@
• Data-constrained law. Validation loss is U-shaped in epochs $e$; the onset of overfitting scales roughly as $e_{\mathrm{opt}}\!\propto\!U_{D}^{0.39}/N^{0.55}$, where $N$ is the model size and $U_{D}$ is the unique data size; e.g., a 10B model on 11T unique tokens tolerates $\sim$1,100 epochs before degradation. We provide practical allocation guidance in Table 3.
@@ 1 Introduction
2025 marks the first year of diffusion language models (DLMs) scaling. Based on the great efforts that laid the theoretical foundation for DLMs (Lou et al., 2023; Shi et al., 2024; Ou et al., 2024; Sahoo et al., 2024), Nie et al. (2025) successfully trained the first large diffusion language model from scratch, competitive to state-of-the-art open-source autoregressive (AR) models (Dubey et al., 2024). Meanwhile, several commercial DLMs emerged, exhibiting superior coding and math performance with remarkably low generation latency at the same time (Google DeepMind, 2025; Khanna et al., 2025; Song et al., 2025). Thereafter, Ni and team (2025) showed that DLMs exhibit much better data learning potential than AR models when data is the bottleneck, a.k.a. "intelligence crossovers", demonstrating a core advantage over AR models under token crisis (Xue et al., 2023; Muennighoff et al., 2023).
@@
DLMs exhibit several modeling advantages over AR models. Their bidirectional attention and diffusion objective enable any-order modeling, allowing data to be modeled in arbitrary directions during both training and inference. This property is particularly beneficial for tasks requiring non-causal dependencies and back-and-forth reasoning, such as coding (Xie et al., 2025; Wu et al., 2025a), mathematics (Google DeepMind, 2025), report generation (Han et al., 2025), etc. DLMs’ bidirectional attention natively support on-the-fly context modification as new content is generated, a desirable feature in these tasks. Multi-token generation is also natively supported by DLMs, providing the foundation for their bleeding fast decoding. Moreover, DLMs spend more parallelable FLOPs at both the inference and training time, leading to their superior data learning capability and potentially stronger reasoning capabilities.
@@
However, the knowledge on how to train large DLMs from scratch is still near blank. Existing studies are largely heuristic or simply extrapolate conclusions from AR models (Nie et al., 2025; Ye et al., 2025). In practice, two scaling laws are of primary interest: (1) the compute-constrained (or compute-optimal) scaling law (Hoffmann et al., 2022), where compute is fixed while model and dataset size are unconstrained; and (2) the data-constrained scaling law (Muennighoff et al., 2023), where dataset size is fixed while model size and compute are unbounded. Both regimes raise key questions about scaling behavior under these restrictions and, more critically, how to optimally allocate the remaining degrees of freedom. Moreover, beyond the classic trade-offs among data, parameters, and compute, additional modeling and optimization choices can also substantially affect the end-of-training performance of language models.
@@
In this work, we will empirically investigate the dependence of language modeling loss and downstream evaluations on all of these factors. We introduce Quokka, the first systematic scaling law for DLMs, covering both the compute-constrained and data-constrained regimes, and studying the key modeling and optimization designs. Specifically, the key contributions we made include:
@@ 2.1 Chinchilla Scaling Law and Its Data-Constrained Version for AR Models
Hoffmann et al. (2022) studies compute-constrained (or compute-optimal) AR pre-training by triangulating evidence from three complementary approaches: (i) Fixed-Parameters: vary training tokens $D$ while holding model size $N$ fixed; (ii) Fixed-FLOPs (IsoFLOP): keep total training compute $C$ fixed while co-varying $N$ and $D$; (iii) Parametric Fit: fit a two-factor loss surface $L(N,D)$ and derive the compute-optimal allocation. Its core parametric law is
@@
When unique data is limited, repeated tokens and excess parameters have diminishing marginal value. Muennighoff et al. (2023) capture this by replacing the raw $(N,D)$ in Equation (4) by their effective counterparts $(N^{\prime},D^{\prime})$:
@@ 2.2 Masked Diffusion Language Models
Why masked diffusion? DLMs adopt a noising–denoising framework over sequences. Among their variants, masked diffusion—also known as absorbing discrete diffusion, which relies on an absorbing transition kernel—has emerged as the most effective formulation (Amin et al., 2025). It preserves discreteness, supports any-order modeling, enables exact position-wise factorization during corruption, and allows flexible likelihood estimation and natively support multi-token prediction. These properties make masked diffusion a strong competitor to AR modeling while retaining many of its practical advantages. Moreover, Ni and team (2025) demonstrate that masked DLMs consistently outperform AR models under data-constrained regimes through more repetitions on data. This advantage is likely rooted in DLMs’ any-order modeling, high compute-parameter ratio, and inherent data augmentation.
@@
i.e., already-revealed tokens stay fixed; masked tokens either remain masked with probability $\frac{1-\alpha_{s}}{1-\alpha_{t}}$ or are revealed by sampling from a data-prediction distribution $q_{0|t}(\cdot\mid x_{t})$ with probability $\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$. A key time-agnostic property (Ou et al., 2024) of masked diffusion is that
@@ 3 Compute-Constrained Scaling Law for Diffusion Language Models
Constrained compute in model training is inevitable—every player in the AGI race faces limited compute budgets while having effectively unlimited model variants to explore. We therefore ask: Given a fixed FLOPs budget, how should one optimally trade off model size against the number of training tokens? Following Hoffmann et al. (2022), we model the DLM training loss, model size, and dataset size using power-law relationships under the limited-compute, infinite-data regime, where each model is trained for a single epoch.
We present two approaches to address this question. First, we conduct extensive IsoFLOPs runs across a range of compute budgets, varying model sizes up to 11B parameters and dataset sizes up to 260B tokens. This allows us to trace the efficient frontier for compute-optimal allocation between model size and dataset size. Second, we fit the power-law loss function to the final training losses obtained from these IsoFLOPs runs. Both approaches converge on the same conclusion: model size and dataset size should scale proportionally with training compute, i.e., doubling $N$ requires doubling $D$, consistent with findings for AR models. However, both approaches also suggest a substantially higher fixed data allocation—roughly $2\text{--}5\times$ that of AR models—for a given FLOPs budget, implying that DLMs are more data-hungry when trained for only a single epoch. Note that Ni and team (2025) shows that DLMs achieve higher data potential under multi-epoch training.
@@ 3.1 Approach 1: IsoFLOPs Profiles
An instructive head-to-head comparison is that the only dense diffusion language model trained from scratch, LLaDA (Nie et al., 2025), which consumed $1.1\times 10^{23}$ FLOPs, adopted a suboptimal parameter–data allocation. As shown in Figure 1 (center, right), the compute-optimal allocation at this budget would be a 15B-parameter model trained on 1.2T tokens, rather than the 8B model with 2.3T tokens they used. We provide a direct comparison between AR models and DLMs under compute constraints in §3.3.
@@ 3.2 Approach 2: Fitting a Parametric Loss Function
The second approach models final training loss as a parametric function of model size $N$ (parameter count) and dataset size $D$. Following Hoffmann et al. (2022), we adopt a functional form based on classical risk decomposition, expressing the loss $L(N,D)$ as:
@@
To estimate the five free parameters $(A,B,E,\alpha,\beta)$, we regress the functional form against our experimental results. Concretely, we minimize the Huber loss (Huber, 1992) between predicted and observed log-losses using the L-BFGS algorithm (Nocedal, 1980):
@@ 3.3 Optimal Model Scaling
Note that Hoffmann et al. (2022) employed three fitting methods. We merge their approaches 1 and 2 into Quokka approach 1, as they are effectively equivalent. Their approach 3, in contrast, reported negative curvature in the $N\rightarrow N_{\text{opt}}$ frontier, yielding lower $N_{\text{opt}}$ estimates. Accordingly, for coefficients other than the irreducible loss $E$, we compare against Chinchilla approaches 1 and 2. The irreducible loss is reported only under their approach 3, i.e., the parametric fit.
@@ 4 Data-Constrained Scaling Law for Diffusion Language Models
In the long run, compute will not be the bottleneck in the pursuit of greater intelligence. According to Common Crawl’s official statistics (Common Crawl, 2025), web data grows roughly linearly, whereas compute for training AI models grows exponentially (Sevilla and Roldán, 2024). Since compute can be scaled both by increasing chip counts and by extending training time, it is effectively unbounded. By contrast, data constitutes the true limiting factor. In particular, certain domains face acute scarcity, including non-English language data, high-quality code, mathematical text, medical data, etc.
@@ 4.1 An Effort in Modeling the Validation Loss with Overfitting
Modeling validation loss is substantially more challenging than modeling pre-training loss. Muennighoff et al. (2023) proposed Equation (7) to capture validation loss, introducing the notion of diminishing "effective model size" and "effective data size," which reflects the intuition that repeated exposure to the same data yields diminishing performance gains. However, this formulation has a critical flaw: it produces a monotonically non-increasing validation loss, which contradicts reality. In practice, repeated training on the same data inevitably leads to increased validation loss due to overfitting, a direct consequence of the bias–variance tradeoff.
@@ 5 Key Modeling and Optimization Choices
We report results on HellaSwag and MMLU, chosen for their popularity and stability across model configurations (Liu et al., 2023; Muennighoff et al., 2024). Their broad adoption allows direct comparison with prior work, making them reliable indicators for assessing the impact of our ablations.
@@
Masked diffusion is often easier to model, as the task reduces to filling in masked positions rather than distinguishing noise from clean tokens. Both theoretical and empirical studies suggest that masked diffusion models generally outperform uniform ones (Amin et al., 2025; Lou et al., 2023). However, direct large-scale comparisons under LLM evaluation settings remain absent.
@@ 5.2 Diffusion Schedules
We examine three common schedules: linear, poly2, and cosine, defined as $\alpha=1-t$, $\alpha=1-t^{2}$, and $\alpha=1-\cos\!\left(\tfrac{\pi}{2}(1-t)\right)$, respectively. Figure 6 (bottom) illustrates their shapes. Cosine assigns higher-than-average probability to masking, poly2 lower, and linear lies in between. From pre-training and evaluation results (Figure 6), cosine performs worst across metrics, while linear consistently outperforms the others in both train/val loss and MMLU. Linear also exhibits lower variance than nonlinear schedules, consistent with Shi et al. (2024). The poly2 schedule achieved better performance on HellaSwag.
@@
A natural intuition in training DLMs is to begin with cleaner data and gradually increase noise, aiming for stronger end-of-training performance (Zhu et al., 2025b). This is straightforward to implement by adjusting the sampling of $t$. In our setup, we use a moving Gaussian window to bias $t$ toward lower values early in training, so the model first learns easier prediction tasks before progressively transitioning to harder ones as the Gaussian window shifts from 0 to 1. Results show that this schedule yields faster loss reduction in the early stages, followed by rising loss as noisier samples dominate. It achieves slightly better end-of-training performance across both benchmarks, suggesting this direction merits further study.
@@ 5.3 Diffusion Loss Formula
Generative masked language models can be trained using either the principled diffusion loss (Shi et al., 2024) or the masked loss (Chang et al., 2022). The diffusion loss is generally regarded as more faithful, since it optimizes a likelihood lower bound and is expected to yield better results. We compare masked generative models trained with diffusion loss (Equation (9)) and MaskGIT loss (Equation (9) without reweighting) over 300B tokens. Surprisingly, despite not optimizing a principled ELBO, MaskGIT achieves consistently comparable performance throughout training and even converges faster on both evaluations. While diffusion loss ultimately delivers stronger end-of-training performance on both benchmarks, this finding highlights the need for further study of how theoretical bounds influence training dynamics.
@@ 5.4 Batch Size and Learning Rate Transferability
Training hyperparameters such as batch size are critical for stability and performance. AR models have well-established scaling laws for these settings (Li et al., 2025). Batch size is closely tied to dataset size, and diffusion language models (DLMs) effectively augment data through noise injection (Ni and team, 2025), often exhibiting higher variance in pre-training loss. Larger batches can mitigate both issues, raising the question of whether DLMs favor larger batch sizes than AR counterparts. Surprisingly, as shown in Figure 9, training dynamics for AR and DLMs are similar: batch size 4096 lags behind 256 and 1024, with the latter two performing comparably. This suggests that changing the training objective does not alter the optimal batch size when data and model architecture are fixed, implying that established AR scaling laws might be able to transfer directly to DLMs.

### v2

@@ Abstract
We introduce Quokka, the first large-scale scaling law for diffusion language models (DLMs), encompassing both compute-constrained and data-constrained regimes, and studying the key modeling and optimization designs. Quokka is a good friend of Chinchilla and provides wider scopes. We hope the results would bring short-term practical guidance in DLMs training and long-term inspirations for the whole AI community. We summarize some takeaways below:
@@
• Data-constrained scaling law. Validation loss is U-shaped in epochs $e$; the onset of overfitting scales roughly as $e_{\mathrm{opt}}\!\propto\!U_{D}^{0.39}/N^{0.55}$, where $N$ is the model size and $U_{D}$ is the unique data size; e.g., a 10B model on 11T unique tokens tolerates $\sim$1,100 epochs before degradation. We provide practical allocation guidance in Table 3.
@@ 1 Introduction
2025 marks the first year of diffusion language models (DLMs) scaling. Based on the great efforts that laid the theoretical foundation for DLMs (lou2023discrete; shi2024simplified; ou2024your; sahoo2024simple), nie2025large successfully trained the first large diffusion language model from scratch, competitive to state-of-the-art open-source autoregressive (AR) models (dubey2024llama). Meanwhile, several commercial DLMs emerged, exhibiting superior coding and math performance with remarkably low generation latency at the same time (deepmind2025geminiDiffusion; khanna2025mercury; song2025seed). Thereafter, ni2025difflm showed that DLMs exhibit much better data learning potential than AR models when data is the bottleneck, a.k.a. "intelligence crossovers", demonstrating a core advantage over AR models under token crisis (xue2023repeat; muennighoff2023scaling).
@@
DLMs exhibit several modeling advantages over AR models. Their bidirectional attention and diffusion objective enable any-order modeling, allowing data to be modeled in arbitrary directions during both training and inference. This property is particularly beneficial for tasks requiring non-causal dependencies and back-and-forth reasoning, such as coding (xie2025dream; wu2025fast), mathematics (deepmind2025geminiDiffusion), report generation (han2025deep), etc. DLMs’ bidirectional attention natively support on-the-fly context modification as new content is generated, a desirable feature in these tasks. Multi-token generation is also natively supported by DLMs, providing the foundation for their bleeding fast decoding. Moreover, DLMs spend more parallelable FLOPs at both the inference and training time, leading to their superior data learning capability and potentially stronger reasoning capabilities.
@@
However, the knowledge on how to train large DLMs from scratch is still near blank. Existing studies are largely heuristic or simply extrapolate conclusions from AR models (nie2025large; ye2025dream). In practice, two scaling laws are of primary interest: (1) the compute-constrained (or compute-optimal) scaling law (hoffmann2022training), where compute is fixed while model and dataset size are unconstrained; and (2) the data-constrained scaling law (muennighoff2023scaling), where dataset size is fixed while model size and compute are unbounded. Both regimes raise key questions about scaling behavior under these restrictions and, more critically, how to optimally allocate the remaining degrees of freedom. Moreover, beyond the classic trade-offs among data, parameters, and compute, additional modeling and optimization choices can also substantially affect the end-of-training performance of language models.
@@
In this work, we will empirically investigate the dependence of language modeling loss and downstream evaluations on all of these factors. We introduce Quokka, the first large-scale scaling law for DLMs, covering both the compute-constrained and data-constrained regimes, and studying the key modeling and optimization designs. Specifically, the key contributions we made include:
@@ 2.1 Chinchilla Scaling Law and Its Data-Constrained Version for AR Models
hoffmann2022training studies compute-constrained (or compute-optimal) AR pre-training by triangulating evidence from three complementary approaches: (i) Fixed-Parameters: vary training tokens $D$ while holding model size $N$ fixed; (ii) Fixed-FLOPs (IsoFLOP): keep total training compute $C$ fixed while co-varying $N$ and $D$; (iii) Parametric Fit: fit a two-factor loss surface $L(N,D)$ and derive the compute-optimal allocation. Its core parametric law is
@@
When unique data is limited, repeated tokens and excess parameters have diminishing marginal value. muennighoff2023scaling capture this by replacing the raw $(N,D)$ in Equation (4) by their effective counterparts $(N^{\prime},D^{\prime})$:
@@ 2.2 Masked Diffusion Language Models
Why masked diffusion? DLMs adopt a noising–denoising framework over sequences. Among their variants, masked diffusion—also known as absorbing discrete diffusion, which relies on an absorbing transition kernel—has emerged as the most effective formulation (amin2025masking). It preserves discreteness, supports any-order modeling, enables exact position-wise factorization during corruption, and allows flexible likelihood estimation and natively support multi-token prediction. These properties make masked diffusion a strong competitor to AR modeling while retaining many of its practical advantages. Moreover, ni2025difflm demonstrate that masked DLMs consistently outperform AR models under data-constrained regimes through more repetitions on data. This advantage is likely rooted in DLMs’ any-order modeling, high compute-parameter ratio, and inherent data augmentation.
@@
i.e., already-revealed tokens stay fixed; masked tokens either remain masked with probability $\frac{1-\alpha_{s}}{1-\alpha_{t}}$ or are revealed by sampling from a data-prediction distribution $q_{0|t}(\cdot\mid x_{t})$ with probability $\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$. A key time-agnostic property (ou2024your) of masked diffusion is that
@@ 3 Compute-Constrained Scaling Law for Diffusion Language Models
Constrained compute in model training is inevitable—every player in the AGI race faces limited compute budgets while having effectively unlimited model variants to explore. We therefore ask: Given a fixed FLOPs budget, how should one optimally trade off model size against the number of training tokens? Following hoffmann2022training, we model the DLM training loss, model size, and dataset size using power-law relationships under the limited-compute, infinite-data regime, where each model is trained for a single epoch.
We present two approaches to address this question. First, we conduct extensive IsoFLOPs runs across a range of compute budgets, varying model sizes up to 11B parameters and dataset sizes up to 260B tokens. This allows us to trace the efficient frontier for compute-optimal allocation between model size and dataset size. Second, we fit the power-law loss function to the final training losses obtained from these IsoFLOPs runs. Both approaches converge on the same conclusion: model size and dataset size should scale proportionally with training compute, i.e., doubling $N$ requires doubling $D$, consistent with findings for AR models. However, both approaches also suggest a substantially higher fixed data allocation—roughly $2\text{--}5\times$ that of AR models—for a given FLOPs budget, implying that DLMs are more data-hungry when trained for only a single epoch. Note that ni2025difflm shows that DLMs achieve higher data potential under multi-epoch training.
@@ 3.1 Approach 1: IsoFLOPs Profiles
An instructive head-to-head comparison is that the only dense diffusion language model trained from scratch, LLaDA (nie2025large), which consumed $1.1\times 10^{23}$ FLOPs, adopted a suboptimal parameter–data allocation. As shown in Figure 2 (center, right), the compute-optimal allocation at this budget would be a 15B-parameter model trained on 1.2T tokens, rather than the 8B model with 2.3T tokens they used. We provide a direct comparison between AR models and DLMs under compute constraints in §3.3.
@@ 3.2 Approach 2: Fitting a Parametric Loss Function
The second approach models final training loss as a parametric function of model size $N$ (parameter count) and dataset size $D$. Following hoffmann2022training, we adopt a functional form based on classical risk decomposition, expressing the loss $L(N,D)$ as:
@@
To estimate the five free parameters $(A,B,E,\alpha,\beta)$, we regress the functional form against our experimental results. Concretely, we minimize the Huber loss (huber1992robust) between predicted and observed log-losses using the L-BFGS algorithm (nocedal1980updating):
@@ 3.3 Optimal Model Scaling
Note that hoffmann2022training employed three fitting methods. We merge their approaches 1 and 2 into Quokka approach 1, as they are effectively equivalent. Their approach 3, in contrast, reported negative curvature in the $N\rightarrow N_{\text{opt}}$ frontier, yielding lower $N_{\text{opt}}$ estimates. Accordingly, for coefficients other than the irreducible loss $E$, we compare against Chinchilla approaches 1 and 2. The irreducible loss is reported only under their approach 3, i.e., the parametric fit.
@@ 4 Data-Constrained Scaling Law for Diffusion Language Models
In the long run, compute will not be the bottleneck in the pursuit of greater intelligence. According to Common Crawl’s official statistics (CommonCrawl2025CrawlSize), web data grows roughly linearly, whereas compute for training AI models grows exponentially (SevillaRoldan2024ComputeGrowth). Since compute can be scaled both by increasing chip counts and by extending training time, it is effectively unbounded. By contrast, data constitutes the true limiting factor. In particular, certain domains face acute scarcity, including non-English language data, high-quality code, mathematical text, medical data, etc.
@@ 4.1 An Effort in Modeling the Validation Loss with Overfitting
Modeling validation loss is substantially more challenging than modeling pre-training loss. muennighoff2023scaling proposed Equation (7) to capture validation loss, introducing the notion of diminishing "effective model size" and "effective data size," which reflects the intuition that repeated exposure to the same data yields diminishing performance gains. However, this formulation has a critical flaw: it produces a monotonically non-increasing validation loss, which contradicts reality. In practice, repeated training on the same data inevitably leads to increased validation loss due to overfitting, a direct consequence of the bias–variance tradeoff.
@@ 5 Key Modeling and Optimization Choices
We report benchmark results on HellaSwag (commonsense reasoning) and MMLU (knowledge), chosen for their popularity and stability across model configurations (liu2023llm360; muennighoff2024olmoe). Their broad adoption allows direct comparison with prior work, making them reliable indicators for assessing the impact of our ablations.
@@
Masked diffusion is often easier to model, as the task reduces to filling in masked positions rather than distinguishing noise from clean tokens. Both theoretical and empirical studies suggest that masked diffusion models generally outperform uniform ones (amin2025masking; lou2023discrete). However, direct large-scale comparisons under LLM evaluation settings remain absent.
@@ 5.2 Diffusion Schedules
We examine three common schedules: linear, poly2, and cosine, defined as $\alpha=1-t$, $\alpha=1-t^{2}$, and $\alpha=1-\cos\!\left(\tfrac{\pi}{2}(1-t)\right)$, respectively. Figure 7 (bottom) illustrates their shapes. Cosine assigns higher-than-average probability to masking, poly2 lower, and linear lies in between. From pre-training and evaluation results (Figure 7), cosine performs worst across metrics, while linear consistently outperforms the others in both train/val loss and MMLU. Linear also exhibits lower variance than nonlinear schedules, consistent with shi2024simplified. The poly2 schedule achieved better performance on HellaSwag.
@@
A natural intuition in training DLMs is to begin with cleaner data and gradually increase noise, aiming for stronger end-of-training performance (zhu2025skyladder). This is straightforward to implement by adjusting the sampling of $t$. In our setup, we use a moving Gaussian window to bias $t$ toward lower values early in training, so the model first learns easier prediction tasks before progressively transitioning to harder ones as the Gaussian window shifts from 0 to 1. Results show that this schedule yields faster loss reduction in the early stages, followed by rising loss as noisier samples dominate. It achieves slightly better end-of-training performance across both benchmarks, suggesting this direction merits further study.
@@ 5.3 Diffusion Loss Formula
Generative masked language models can be trained using either the principled diffusion loss (shi2024simplified) or the masked loss (chang2022maskgit). The diffusion loss is generally regarded as more faithful, since it optimizes a likelihood lower bound and is expected to yield better results. We compare masked generative models trained with diffusion loss (Equation (9)) and MaskGIT loss (Equation (9) without reweighting) over 300B tokens. Surprisingly, despite not optimizing a principled ELBO, MaskGIT achieves consistently comparable performance throughout training and even converges faster on both evaluations. While diffusion loss ultimately delivers stronger end-of-training performance on both benchmarks, this finding highlights the need for further study of how theoretical bounds influence training dynamics.
@@ 5.4 Batch Size and Learning Rate Transferability
Training hyperparameters such as batch size are critical for stability and performance. AR models have well-established scaling laws for these settings (li2025predictable). Batch size is closely tied to dataset size, and diffusion language models (DLMs) effectively augment data through noise injection (ni2025difflm), often exhibiting higher variance in pre-training loss. Larger batches can mitigate both issues, raising the question of whether DLMs favor larger batch sizes than AR counterparts. Surprisingly, as shown in Figure 10, training dynamics for AR and DLMs are similar: batch size 4096 lags behind 256 and 1024, with the latter two performing comparably. This suggests that changing the training objective does not alter the optimal batch size when data and model architecture are fixed, implying that established AR scaling laws might be able to transfer directly to DLMs.